### v1

# Predicting the Geothermal Gradient in Colombia: a Machine Learning Approach

Abstract
Accurate determination of the geothermal gradient is critical for assessing the geothermal energy potential of a given region. Of particular interest is the case of Colombia, a country with abundant geothermal resources. A history of active oil and gas exploration and production has left drilled boreholes in different geological settings, providing direct measurements of the geothermal gradient. Unfortunately, large regions of the country where geothermal resources might exist lack such measurements. Indirect geophysical measurements are costly and difficult to perform at regional scales. Computational thermal models could be constructed, but they require very detailed knowledge of the underlying geology and uniform sampling of subsurface temperatures to be well-constrained. We present an alternative approach that leverages recent advances in supervised machine learning and available direct measurements to predict the geothermal gradient in regions where only global-scale geophysical datasets and course geological knowledge are available. We find that a Gradient Boosted Regression Tree algorithm yields optimal predictions and extensively validate the trained model. We show that predictions of our model are within 12% accuracy and that independent measurements performed by other authors agree well with our model. Finnally, we present a geothermal gradient map for Colombia that highlights regions where futher exploration and data collection should be performed.
keywords:
Geothermal Energy, Machine Learning, Colombia, Geothermal Gradient, Prediction
1 Highlights
  • A machine learning approach for geothermal gradient prediction using only global-scale geophysical datasets and low resolution geological maps.
  • Model predictions show good agreement with independent measurements and high accuracy.
  • Model allows for the generation of a new geothermal gradient map of Colombia highlighting the potential of unexplored regions.
  • Model predictions follow regional trends established by prior geological knowledge.
  • Model predictions confirm high geothermal gradients in the Amazon, an unexplored region, giving a baseline to start studying the area.
2 Introduction
The tectonic and geological characteristics of Colombia give rise to thermal anomalies and regions exhibiting elevated geothermal gradients, which represent promising sources of geothermal energy. However, performing accurate estimations of the country’s potential for geothermal energy generation remains challenging. The most current geothermal gradient map for Colombia covers roughly half of the country’s territory (Alfaro et al., 2009), for the rest no systematic determinations of the geothermal gradient are available. Some studies have tried to address this issue. Uyeda and Watanabe (1970) observed normal to subnormal geothermal gradient values across South America, with higher values in the Andes related to geothermal activity. Bachu et al. (1995) focused on the Llanos Basin, finding that geothermal gradients decrease with depth and westward. In northwestern Colombia, Quintero et al. (2019) estimated geothermal gradients using grid systems and aeromagnetic data. Unfortunately, all of these studies are either focused on narrow areas of interest or scales that are too large for meaningful estimations of the country’s geothermal potential. Measurements from boreholes drilled by the oil and gas industry provide the best direct constraint on the geothermal gradient, in the case of Colombia they are either very sparse or clustered to specific regions, as the need to decarbonize the economy grows, it is unlikely that the past pace of oil exploration activity will continue. Therefore, a different approach is needed if Colombia is to realize its clean energy production objectives (Alfaro et al., 2021).
It is seldom impossible to develop comprehensive geothermal gradient models for an entire country using only direct borehole measurements (Burton-Johnson et al., 2020), indirect approaches are necessary. Geophysical methods provide complementary estimates of the geothermal gradient over large spatial scales and depths that are inaccessible by boreholes alone; magnetic and electrical surveys exploit the thermal properties of rocks to estimate temperature gradients. While ideal, due to their cost such surveys are typically performed once a narrow target of interest has been identified. Numerical thermal modeling has emerged as a useful tool for estimating geothermal gradients as it allows for the prediction of the entire temperature field of the region’s subsurface (Békési et al., 2020). Thermal models require detailed knowledge of the underlying geology, direct temperature and heat flow measurements and involve a significant number of assumptions. For Colombia, while in principle modeling could be performed it is unlikely to yield accurate results.
Data-driven approaches have become viable alternatives for the estimation of the earth’s thermal properties. Rezvanbehbahani et al. (2017) used a gradient boosting regression tree algorithm and a catalog of global geological features to produce a new geothermal heat flow (GHF) map of Greenland, their predictions were consistent with regional geology, tectonics and ice core measurements. A similar approach was used to predict GHF in Antarctica by Lösing & Ebbing (2021), it incorporated regional datasets to improve predictions. Matiz-León (2023) emphasized the importance of statistical methods and spatial prediction for estimating geothermal gradients, especially in areas lacking in-situ data. Advances in machine learning and the increasing availability of regional and global geological datasets allow for the development of cost effective approaches to provide initial baselines of the thermal properties of entire countries.
In this work we present a Gradient Boosted Regression Tree algorithm to accurately predict the spatial distribution of geothermal gradients across a region. The approach only requires global scale geological and geophysical datasets, coarse geological knowledge and sparsely distributed borehole temperature measurements to produce accurate results. The paper is organized as follows: In section 2 we describe the details of the algorithm, in section 3 we present the data and the strategy used to train the model, in section 4 we provide extensive validation of our model and present a new map of geothermal gradients for Colombia. The predicted map highlights regions displaying relatively high geothermal gradients, where future exploration surveys should be conducted. Further model validation is provided by the fact that predictions show good agreement with indirect geophysical measurements of the geothermal gradient performed in the Amazon.
3 Method
To achieve precise quantitative predictions within the geothermal context, it is necessary to utilize a regression algorithm that can effectively capture the relationship between the Geothermal Gradient and its geological setting. Our approach utilizes gradient boosting regression, a technique within the domain of supervised machine learning. This method systematically develops an ensemble of regression trees, where each tree is designed to incrementally adjust to the gradient of the loss function from the previous one, following the methodology outlined by Friedman (2001).
Regression trees are particularly adept at identifying complex, nonlinear relationships by partitioning data recursively based on specific criteria, such as maximum depth, number of leaves, and minimum samples per leaf. In the gradient boosting framework, each iteration involves training a new tree on a subset of the data to accurately predict gradient descent steps. By aggregating these trees into an ensemble model, we enhance our capacity to predict outcomes for new data samples. This method stands out for its straightforward interpretability and operational simplicity, qualities that are often highlighted in the literature (Lösing & Ebbing, 2021).
3.1 Enhanced Gradient Boosting with XGBoost
In this study, we utilize the XGBoost algorithm, a refined version of gradient boosting regression known for its high efficiency and robust regularization techniques. Developed by Chen & Guestrin (2016), XGBoost is particularly adept at handling complex models and large datasets. The methodology involving XGBoost encompasses the following aspects:
  • Objective Function: XGBoost optimizes a comprehensive objective function that integrates both the loss function and regularization to mitigate overfitting and enhance model performance. The objective function is meticulously crafted as follows:
    $\mathcal{L}(\phi)=\sum_{i}l(\hat{y}_{i},y_{i})+\sum_{k}\Omega(f_{k})$ (1)
    where $\mathcal{L}(\phi)$ represents the total loss to be minimized. The first term, $\sum_{i}l(\hat{y}_{i},y_{i})$, quantifies the prediction error through a loss function $l$, contrasting the predicted outcomes $\hat{y}_{i}$ against the actual targets $y_{i}$. The second term, $\sum_{k}\Omega(f_{k})$, introduces regularization, encapsulating the complexity of the model to prevent overfitting. This dual approach ensures a balanced focus on achieving high predictive accuracy while maintaining model simplicity.
  • Regularization Component: The regularization term $\Omega(f_{k})$ plays a pivotal role in XGBoost’s formulation, explicitly designed to control model complexity:
    $\Omega(f_{k})=\gamma T+\frac{1}{2}\lambda\lVert w\rVert^{2}$ (2)
    Here, $T$ signifies the total number of leaves in tree $k$, and $w$ denotes the vector of scores on the leaves. The parameter $\gamma$ serves as the complexity cost assigned to each additional leaf, penalizing the growth of the tree structure. The term $\lambda$ is the L2 regularization parameter on the leaf weights, reducing their magnitude to enhance generalization capabilities. Collectively, these components of $\Omega(f_{k})$ ensure that the model remains both effective and interpretable.
  • Gradient and Hessian Optimization: The optimization strategy in XGBoost employs the Newton-Raphson method, leveraging second-order derivatives for a refined adjustment of the model. The approximation for optimization at iteration $t$ is given by:
    ~$\tilde{\mathcal{L}}(t)=\sum_{i=1}^{n}\left[g_{i}f_{t}(x_{i})+\frac{1}{2}h_{i}f_{t}^{2}(x_{i})\right]+\Omega(f_{t})$ (3)
    In this equation, $g_{i}$ and $h_{i}$ represent the first and second-order gradients of the loss function with respect to the predictions, facilitating a nuanced understanding of the error landscape. This methodological choice allows XGBoost to more accurately pinpoint the direction and magnitude of model updates, markedly improving convergence speed and prediction accuracy.
  • Shrinkage and Column Subsampling: Beyond regularization, XGBoost incorporates shrinkage and column subsampling to further enhance model robustness. Shrinkage, through the application of a learning rate, systematically scales down updates to the model, enabling more gradual learning and reducing the risk of overfitting. Column subsampling introduces randomness into the feature selection process for tree construction, promoting diversity among the trees and contributing to a more generalized model performance.
The implementation of XGBoost in our study uses these components to create a predictive model that is accurate and capable of handling the complexities inherent in geothermal gradient data.
3.2 Model Building and Evaluation
The development and validation phases of the model were structured to enhance its learning efficacy and accuracy in predicting geothermal gradients. Our methodological framework incorporates several critical steps that collectively improve the model’s performance.
The initial phase involved the partitioning of the dataset into training and testing subsets with an 80/20 split, a standardized step for model construction, and subsequent validation with independent test data. This split facilitates the proper evaluation of the model’s ability to generalize. Furthermore, we performed hyperparameter optimization to fine-tune the XGBoost model, enhancing its predictive accuracy.
Following the hyperparameter tuning, the XGBoost model was trained using a squared error loss function. The training process incorporated a five-fold grid search approach to fine-tune parameters such as shrinkage, maximum tree depth, and subsampling rate. This iterative refinement ensured the model’s robustness and adaptability.
The validation of the model used a quantitative framework influenced by Lösing & Ebbing (2021) and Rezvanbehbahani et al. (2017), focusing on absolute and relative error metrics. For the assessment of absolute error in the test set, we computed both the root mean square error (RMSE) and mean absolute error (MAE). Furthermore, normalized root mean square error (nRMSE) and normalized mean absolute error (nMAE) were calculated to provide standard and relative error measurements that are invariant to rescaling, facilitating comparative analysis.
The coefficient of determination, or R-squared ($R^{2}$), was calculated for both the test and training sets. The $R^{2}$ for the test set provides information on how well the predicted geothermal gradients align with the actual observed values, indicating the accuracy of the model in unseen data. Similarly, the $R^{2}$ value for the training set, referred to as the model score, reflects the proportion of variance in the geothermal gradient data that is predictable from the input variables, providing a gauge of the model’s efficacy during the learning phase.
Additionally, the process included analyzing the feature importance for each tree, averaged across the model. This analysis revealed how each feature reduced the uncertainty at its split point, weighted by the observation count in each node. These insights were crucial for improving the prediction model and understanding the factors that affect the predictions of geothermal gradients.
4 Data and features
In this section, we explain the data used for constructing our predictive model. The target feature $y$ was the Apparent Geothermal Gradient (AGG) measurements from over 4000 well logs along the colombian territory, and for the predictor features $X$ we use certain geological and geophysical features which are linked to each AGG value.
4.1 Apparent Geothermal Gradient
Colombian AGG measurements were taken from the Alfaro et al. (2009) Colombian Geothermal Gradients map, which includes the data from over 4000 wells (Figure 0(a). The methodology they adopted for calculating the geothermal gradient for each well involved the following steps:
  1. 1.
    Measurement of Bottom Hole Temperature (BHT): BHT measurements from gas and oil wells were primarily used. These measurements represent the maximum temperature recorded in a well. Due to the drilling process, these measurements are often lower than the actual formation temperature.
  2. 2.
    Temperature Measurement Correction: Various models and methods were used to correct the effects of drilling on formation temperatures. The most widely used method for correcting BHT data was the AAPG method, which compares BHT measurements with equilibrium temperatures in wells in Louisiana and West Texas. The correction was achieved using a third-order polynomial equation, with BHT correction as a function of depth.
  3. 3.
    Data Sources and Treatment: Data were obtained from the Finder and LogDB tables supplied by EPIS. The organisation of data involved forming pairs of depth-temperature data, and the selection process involved criteria for data quality. Units were converted from degrees Fahrenheit and feet to degrees Celsius and kilometres. Temperatures were corrected using the empirical method of AAPG.
  4. 4.
    Estimation of Individual Well Gradients: The maximum gradient within the upper crust was assumed to be vertical and calculated using the equation $G=\Delta T/\Delta P=(TBHT-Ts)/Z$, where $TBHT$ is the corrected bottom hole temperature, $Ts$ is the surface temperature, and $Z$ is the depth. Surface temperature was estimated based on correlation with elevation.
  5. 5.
    Interpolation and Map Creation: A minimum curvature interpolation method was used with specific parameters in the Geosoft software. The map was edited in ArcMap, including official coverages of sedimentary basins and coverage of points with conventions for differentiating the magnitude of the estimated gradient and the measured BHT depth.
The geothermal gradient values are predominantly ranging between 10 and 40 °C/Km, as shown in Figure 0(b). The data exhibit a unimodal distribution, characterized by a pronounced peak and a right-skewed asymmetry, indicating a longer tail of higher geothermal gradient values.
The distribution suggests that we are working with an imbalanced dataset, while the majority of measurements are clustered at lower gradients, a subset of regions exhibits elevated gradients, potentially highlighting areas of heightened geothermal activity. This skewness towards higher values contributes to an increased error rate in the model when predicting in these less represented higher gradient zones.
To deal with underrepresented high-gradient values, we defined a high-gradient threshold of 33.86 °$\textdegree C/km$, which corresponds to the value at the 94th percentile (Figure 0(b)). We then calculated the ratio between the low (below the threshold) and high (above the threshold) gradient values and defined a weight of 8 for samples above the threshold of high gradient, which was adjusted based on model evaluation.
Furthermore, it is important to consider that the geothermal gradient data have an inherent bias: Although the measurements were made mainly in basins, a considerable amount of the data is close to Andean and volcanic areas, so measurements made in areas farther away tend to be underrepresented and affected the prediction for areas where there are no measurements. To correct for this bias, we defined each well located more than 50 km from the nearest volcano as ”nonvolcanic” and assigned a weigh of 4 to those wells.
Finally, we constrained the training and test sets to a specific range: from the 1st percentile value of 14.87 ${}^{\circ}$/km to the 99th percentile value of 43.80 ${}^{\circ}$/km (Figure 0(b)). This selection ensures that the model is trained and evaluated in a confidence range.
Figure 1: (a) Measured geothermal gradients in Colombia, (b) Histogram of apparent geothermal gradient values indicating a right-skewed distribution. Dashed lines indicated the representative percentiles used to train and test the model.
4.2 Geological and Geophysical Data
To predict the apparent geothermal gradient, we utilized a diverse array of geological and geophysical data. This data provided insight into the subsurface conditions necessary for accurate geothermal gradient estimation.
The AGG dataset includes latitude and longitude, for georeferencing well sites. This geospatial positioning is used to integrate the spatial variations of geothermal data into the model. Alongside these coordinates, topography and elevation play a significant role. The topographic elevation data, derived from well data and NASA’s global Earth Digital Elevation Model (DEM) as of 2021, provides a vertical reference point which is essential for geothermal gradient measuerements.
Furthermore, subsurface characteristics are another key aspect of the dataset. Insights into the thermal structure of the lithosphere and the behavior of heat at depth are garnered from the depths to the Moho discontinuity and the Curie isotherm. These are sourced from studies by Camacho et al. (2014) and Uieda & Barbosa (2017). Moho discontinuity and elevation data are strictly correlated to litosphere thickness, which is a great predictor for the goeothermal gradient in the continent (Kolawole & Evenick, 2023); Studies from Quintero et al (2019) and Ponce & Hernández (2014) show that there is an inverse correlation between the thermal gradient and the Curie isotherm.
Additionally, the presence of geophysical anomalies significantly influences the dataset. Magnetic anomalies, documented in the WDMAM V2 map by Dyment et al. (2020), are important as they potentially correlate with AGG. Deepest magnetic sources can be associated to Curie depth (Lösing & Ebbing, Op. Cit).
Gravity anomalies are strong indicators of the crustal composition as well. We use Free Air anomalies and Vertical Gravity Gradients from Sandwell et al. (2021) and Pavlis et al. (2012) datasets, it can indicate variations in the subsurface composition and structure. These variations are essential in understanding the distribution of subsurface heat and are influenced by crustal properties, such as the density variations and their correlation with rock types. The Bouguer anomaly was calculated using the Free Air anomaly with the Bouguer correction, and using this parameter allows us to reduce the elevation influence in the gravitational anomaly. After testing, we decided to keep the three gravity features.
The dataset also considers fault data, specifically the proximity to faults, calculated using data from Gómez et al. (2020) and analyzed through a Nearest Neighbors Algorithm. This proximity provides insights in evaluating potential permeability and fluid flow, which can significantly affect subsurface thermal regimes. Active faults, in particular, identified by Veloza et al. (2012), are of immense importance due to their potential role in localizing geothermal systems.
Lastly, the depth to basement rocks (igneous and metamorphic) is a critical factor in understanding the influence of subsurface rock types on thermal behavior, as they are a main source of radiogenic heat production compared to sedimentary rocks (Hasterok et al., 2018). Due to the absence of a specific basement depth model, an alternative method is employed using the geological map by Gómez et al. (2015). This method involves classifying rock outcrops as either basement or non-basement (sedimentary and deposits), and measuring the distance from each well site to the nearest basement outcrop. This assessment helps in understanding the impact of these rocks on geothermal gradients.
The integration of geospatial coordinates, topographical data, subsurface characteristics, geophysical anomalies, fault data, and the proximity to basement rocks forms a comprehensive framework for accurately modeling and understanding the subsurface geothermal behavior, reflecting the multidisciplinary nature of geothermal research. Each of these factors contributes uniquely to understand the gradient distribution, underlining the complex interplay of geological, geophysical, and spatial factors in geothermal exploration and modeling.
Detailed descriptions of these variables are provided in Table 1 and two graphical examples are shown in figure 2. The proper interpretation and utilization of this data are crucial for the predictive modeling of geothermal gradients across Colombia. The features were directly interpolated to obtain values at specific geographic points without resampling, the dataset with the lowest resolution (Moho Depth) has a 0.4° latitude-longitude grid.
Table 1: Geological and geophysical features used in the study with their respective sources.
Feature Description Source
Elevation (m)
Height above sea level.
NASA Visible Earth (2021)
Moho Depth (m)
Depth to the Mohorovičić discontinuity.
Uieda & Barbosa (2017)
Curie Depth (Km)
Depth at which rocks in a specific geographical area encounter the Curie temperature
Camacho et al. (2014)
Magnetic Anomaly (nT)
Earth’s magnetic field anomaly at the point.
Dyment et al. (2020)
Free Air Anomaly (mGal)
Gravity anomaly after Free Air correction
Sandwell et al. (2021)
Bouguer Anomaly (mGal)
Gravity anomaly after Free Air correction and Bouguer corrections
Sandwell et al. (2021)
Vertical Gravity Gradient (E)
rate of change of vertical gravity ($gz$) with height ($z$)
Pavlis et al. (2012)
Faults (m)
Distance to the nearest normal, reverse or strike-slip fault.
Gómez et al., (2020)
Active Faults (m)
Distance to the nearest active fault.
Veloza et al., (2012)
Distance to basement (m)
Surface distance to igneous or metamorphic outcrops
Gómez et al., (2015)
Volcanic Domain
Boolean variable to define whether the well is located in a volcanic zone or not.
Gómez et al., (2020)
AGG measurements (°C/Km)
Target variable, measured from bottom hole data
Alfaro et al., (2009)
Figure 2: Side-by-side maps of Basement rocks outcrops and Bouguer anomaly with the measured gradient locations.
4.3 Hyperparameter tuning
To optimize the performance of the model, we defined a grid of hyperparameters. The code iterates through these hyperparameter combinations, training the model and evaluating mean squared error (MSE) on the test set for each configuration. The process systematically identifies the combination of hyperparameters that minimizes MSE, ensuring the model is fine-tuned for optimal predictive accuracy on the given dataset (Table 2).
Table 2: Best Hyperparameters after Model Tuning
Hyperparameter Value Meaning
max_depth
7
Maximum depth of a tree.
learning_rate
0.05
Step size shrinkage used in updates.
gamma
0.15
Minimum loss reduction required to make a further partition.
min_child_weight
10
Minimum sum of instance weight needed in a child.
colsample_bytree
1
Fraction of features to be randomly sampled for each tree.
reg_alpha
0.5
L1 regularization term.
reg_lambda
2
L2 regularization term.
subsample
0.7
Fraction of training data to be randomly sampled for each boosting round.
5 Results
The result we present for this study is a predicted geothermal gradient map for all the Colombian territory. The accuracy of the methodology is evaluated using the test subset derived from the initial dataset partitioning. We examine the key performance metrics: MAE, RMSE, nMAE, nRMSE and the coefficient of determination ($R^{2}$) to quantify the predictive accuracy of the model. Additionally, we employ visualizations to facilitate a comprehensive analysis: plots for actual versus predicted values, diagrams illustrating feature importances, and assessments delineating the influence of individual features on the model’s output.
A geospatial comparison of actual data against model predictions is conducted to discern the model’s spatial prediction accuracy. This comparison not only substantiates the model’s validity, but also underscores any region-specific deviations.
5.1 Model Validation
The initial phase of model evaluation involved scrutinizing its performance in both the training and the test subsets, the latter representing previously unseen data. For the training set, the model achieved an MAE of 1.8760, an RMSE of 2.5069, and a coefficient of determination ($R^{2}$) of 0.7670. These metrics are typical for models evaluated on their training data. Conversely, the test data got an MAE of 2.6130, an RMSE of 3.6430, and an $R^{2}$ of 0.5156.
The disparity in $R^{2}$ could be attributed to several factors, including the underrepresentation of higher geothermal gradient values, the geographic bias of the measurements and the inherently complex nature of the dataset. The test set’s normalized error metrics, nMAE and nRMSE, were recorded at 0.06 and 0.12 respectively, indicating that our predictions are expected to be within a range from 6% to 12% of the correct values.
A detailed examination of the residuals, as depicted in Figure 3, reveals a random distribution around the zero line. This pattern suggests an absence of bias in the residuals, corroborated by a residual mean close to zero, indicating that the model is on average neither overestimating nor underestimating significantly. However, points with high residuals highlight instances where the model performance was suboptimal, though most residuals cluster near zero.
Figure 3: Residuals distribution of the model.
Regarding the plot of actual versus predicted values, illustrated in Figure 4, the concentration of points along the ideal prediction line indicates proficient modeling of the target values. The dispersion of points as the target value increases suggests increasing variance in the model’s predictions. The model demonstrates higher accuracy for lower actual values while tending to underestimate higher actual values. The overlap of training (blue) and test (red) data points in the plot is indicative of good model generalization, evidenced by the similarity in performance across the training and test datasets. This overlap is a positive sign, suggesting that the model is not overfitting to the training data and most of the variance in the test set is towards the high gradient values.
Figure 4: Actual vs. Predicted values plot comparing the training and test subsets.
Evaluating the model’s geospatial predictive performance is crucial for understanding its applicability in practical scenarios. This involves a comparative analysis of observed geothermal gradient values against the model’s predictions.
Additionally, a side-by-side visualization of the actual and predicted geothermal gradient values was carried out, as shown in Figure 5. The actual values map is a representation of empirically measured geothermal gradients used for the training , while the predicted values map is derived from the model estimations. The consistency of geothermal gradients across both maps serves as an indicator of the model’s spatial accuracy.
Figure 5: Side-by-side comparison of actual and model-predicted geothermal gradient values, highlighting the model’s spatial prediction accuracy.
The map also shows regions where the model’s predictions deviate from measured data (as it’s evidenced in the Figure 4), those discrepancies suggest potential areas for model enhancement or may pinpoint inherent limitations within the data or the model’s structural assumptions. The tendency of the model to underestimate geothermal gradients, specially when they are higher, indicates possible gaps in its representation of heat source distribution or heat transfer mechanisms, and may require data enhancement.
Comparing actual with predicted values yields critical insights into the model’s predictive strengths and weaknesses. Such an assessment is imperative to refine the model’s predictive capabilities, thereby enhancing its contribution to geothermal exploration and resource evaluation.
5.2 Feature importance analysis
Feature importance provide insight into which variables are most informative for making accurate predictions. Multiple techniques were utilized to determine the importance of the features, each offering a unique perspective on the data.
Weight-Based Importance
The weight-based feature importance, as shown in Figure 6, indicates that elevation and Moho depth are the most frequently used features in the model to make decisions; however, the importance is equilibrated. This highlights the model’s reliance on all the data to predict the target variable, especially topography, subsurface discontinuities, and structural influence.
Figure 6: Weight-based feature importance plot.
Gain-Based Importance
Figure 7 presents the gain-based feature importance, which reflects the improvement in accuracy brought by a feature to the branches, how the features contributed to minimize the loss function and, subsequently, to the model’s performance. Basement proximity in surface emerges as the most significant predictor, however, most of the features importances are equilibrated.
Figure 7: Gain-based feature importance plot.
Mutual Information
Mutual information scores, as shown in Figure 8, quantify the dependency between each feature and the target variable. Moho depth is seen to be highly informative, along with the geophysical and structural data.
Figure 8: Mutual Information with the target variable.
We also evaluated each individual feature using the model score, the R2 score, the normalized root mean square error (nRMSE), and the normalized mean absolute error (nMAE). As shown in Figure 9, each metric is visualized for every feature after training the model with each individually.
Moho depth and geophysical measurements, such as the Free Air Anomaly, consistently demonstrated high importance as isolated predictors. In contrast, the topography and Curie depth showed lower individual relevance.
Figure 9: Feature importance evaluated by various metrics for each individual feature. Values are normalized.
Subsequently, the model performance was assessed as a function of the number of features used (Starting with the less important according to Feature Importance), as depicted in Figure 10. The impact of each feature on the training set (model score) and the test set ($R^{2}$ score) increases with the number of features, indicating improved precision and variance explanation, respectively. However, these scores flatten after including approximately nine features, suggesting that additional features would not contribute significantly to the predictive capability of the model regarding error metrics. However, it is important to recognize that each feature plays a crucial role in enhancing the model’s spatial accuracy, and omitting any feature could therefore significantly affect the outcomes.
Figure 10: Model performance as a function of the number of features included, starting from the less important (gain-based).
The analyses suggest that a subset of features provides the most predictive power without introducing unnecessary complexity or risking overfitting. This balance is crucial for developing a model that is both accurate and generalizable.
While some features have significant predictive power, the model reaches a performance threshold beyond which additional features do not yield further benefits. This information is important for refining the model to focus on the most impactful features and avoid overfitting by excluding less informative ones.
5.3 Spatial Distribution of Predicted Geothermal Gradients
The spatial distribution of the predicted geothermal gradients is a key factor in assessing the utility of the model for geothermal exploration. Figure 11 illustrates the predicted geothermal gradients in the study area. The color gradient from blue to red indicates increasing values, with warmer colors representing higher geothermal potential.
Figure 11: Predicted geothermal gradient map showing the spatial distribution of geothermal potential across the region.
The map offers practical information on areas with significant geothermal activity, marking them as prime candidates for detailed exploration and potential development. The effectiveness of the predictive model lies in its ability to identify these spatial variations and key trends. This characteristic makes it a valuable asset in guiding exploration strategies towards the most promising geothermal energy locations.
5.3.1 Analysis of Predicted Gradients
The model’s output indicates a distinct sensitivity to geothermal indicators, revealing high-gradient values that are suggestive of underlying geothermal resources. These values are particularly pronounced in regions characterized by volcanic activity or where the crust is considerably thinner.
A standout feature of the model is its predictive accuracy in the Amazon basin, where data scarcity has previously been a barrier. The model’s gradient predictions for this area not only fill a critical data gap but also exhibit remarkable agreement with the findings from the Pimentel & Hamza (2012) study of Brazil’s portion of the basin. This concordance is a testament to the model’s robustness and its ability to extrapolate from known data to regions lacking in situ measurements.
In terms of new discoveries, the model highlights areas with potentially high geothermal gradients that are not documented in existing gradient maps. These predicted hotspots, particularly in the uncharted Amazon region, present opportunities for exploration and possible development of geothermal resources.
The model has proven to be a valuable asset in geothermal exploration, especially for the Amazon of Colombia, where it predicts high geothermal gradients in the absence of empirical data. Its alignment with the established geothermal gradients from Pimentel & Hamza’s work further validates its predictive capacity. As the model continues to be refined, its role in the identification and assessment of geothermal resources is expected to become critical, supporting a more informed and strategic approach to energy exploration in the region.
5.3.2 Implications for Geothermal Exploration
The spatial distribution map and the histogram of the predicted geothermal gradients provide critical guidance for geothermal exploration. These tools enable the strategic prioritization of regions for further investigation, allowing for a focused allocation of resources to areas with the highest geothermal potential.
The regions with the highest predicted gradients warrant immediate attention for exploration and data acquisition. These areas, flagged by the model, are likely candidates for sustainable geothermal development. In contrast, areas with lower predicted gradients can be investigated for baseline data collection, contributing to a more comprehensive understanding of the geothermal profile of the region.
Our predictive model can play a pivotal role in the early stages of exploration, for example, inform decision-making processes regarding site selection for geothermal power plants and drilling operations. By leveraging the model’s predictions, stakeholders can enhance their exploration strategies, targeting locations with a higher likelihood of geothermal activity, and thereby advancing the exploration for renewable energy sources. This optimized approach promises not only greater efficiency in resource utilization but also a higher success rate in developing geothermal energy projects.
6 Discussion
The evaluation of our predictive model for the estimation of geothermal gradients has been extensive, incorporating a variety of metrics and visual analyzes. Training set metrics, MAE, RMSE, and R-squared, along with the corresponding test set metrics provide a comprehensive view of the model performance.
The model shows strong performance in the training set, with an MAE of 1.87, an RMSE of 2.50, and an R-squared value of 0.77, as indicated in Section 5.1. These values reflect a high degree of accuracy and a strong correlation between the predicted and actual geothermal gradients, suggesting that the model is capable of capturing the underlying patterns within the training data. However, when applied to the test set, the model’s performance drops, with an MAE of 2.61, an RMSE of 3.64, and an R-squared value of 0.51, as detailed in Section 5.1. The observed increase in error and the decrease in the $R^{2}$ value on the test set, as opposed to the training set, may suggest possible overfitting, particularly in the context of estimating higher values of geothermal gradients where the data are less abundant. Despite this, the model’s ability to generalize is demonstrated to be effective, as illustrated in Figure 4.
Comparative analysis with related studies highlights the effectiveness of our model. For example, the study by Lösing & Ebbing (Op. Cit) in Antarctica achieved an nRMSE of 0.29 and an of $R^{2}$ 0.45, while Rezvanbehbahani et al. (OP. Cit) in Greenland recorded an nRMSE of 0.15 and an $R^{2}$ of 0.6. Our results, with an nRMSE of 0.12 and an $R^{2}$ of 0.52, not only reflect a lower relative error compared to these studies, but also show a balanced performance in terms of the determination coefficient. This comparison highlights our model’s accuracy and reliability across different geological settings.
The geospatial distribution of the predicted geothermal gradients, illustrated in Figure 11, highlights the spatial prediction capabilities of the model. The alignment of high-gradient regions in the actual vs. predicted maps (Figure 5) and the regions of significant discrepancy provide a nuanced understanding of the spatial accuracy of the model.
It is important to note that our model effectively predicted geothermal gradients in data-limited areas like south-western Colombia, including the Amazon and eastern Llanos, that were consistent with Pimentel & Hamza’s (2012) study, which observed higher gradients in the north-western Amazon basin. Training data shows an isolated low-gradient data point near the Amazon (Figure 5), however, that point has not influenced model training or prediction and can therefore be considered an outlier given the geological context of the south-west Colombia.
Our feature importance analyses, comprising weight-based, gain-based, and mutual information methods, showcased the varying impact of different features. Elevation and basement proximity consistently emerged as the main predictors in multiple importance measures, as discussed in Sections 5.2. Moho depth is also one of the most important predictors, it´s relation with the lithospheric thickness suggests an agreement with the correlation of this tectonic feature with the thermal gradient defined by Kolawole & Evenick (2023). The geophysical properties, such as Bouguer or Free Air Anomaly, have been highlighted as significant by certain importance methods but appear less influential in others.
The observed discrepancies between training and test performance may indicate the need of further model refinement or data enhancement. Strategies such as feature engineering, parameter optimization and integration of more specific data such as depth to basement or lithospere thickness could be employed to enhance the model’s predictive power and generalizability.
7 Conclusion
This study has successfully developed a machine learning model to estimate geothermal gradients across Colombia, marking a significant contribution in the field of geothermal energy exploration. By utilizing diverse geophysical and geological data, and gradient boosting regression; we have created a model that displays accurate gradient predictions across unexplored regions in Colombia.
Our model illustrates the potential of machine learning in enhancing the understanding of geothermal gradients distribution, specially within the context of Colombia, where there are vast regions whose geothermal energy has been underexplored (or not explored at all) due to the lack of extensive well data.
The incorporation of a wide range of geological and geophysical features was crucial to our approach, which allowed for a better understanding of the factors influencing geothermal gradients and heat distribution within the subsurface. The findings of this study mark a substantial step forward in geothermal exploration efforts, providing a promising tool for the identification of regions with significant geothermal potential. This advancement is expected to greatly assist in the prioritization of future geothermal exploration plans, making it a valuable asset for both researchers and industry professionals.
As we look to the future, the study encourages further research to refine and enhance the model’s accuracy. This includes the integration of local or regional data instead of global data, and the exploration of additional features that may influence geothermal gradients. Finally, we highlight the capability of Machine learning to contribute, through continued improvements, to the sustainable development of Colombia’s geothermal resources and other renewable energy sources.
8 Data Availability
The data supporting the findings of this study are openly available in the GitHub repository associated with this project. Interested parties can access the dataset, additional resources, and the code used for the analyses at the following URL:
This repository includes all relevant data, scripts, and additional materials required to replicate the study’s results or to further explore the geothermal gradient estimation model. By providing the data in a public domain, we encourage transparency, reproducibility, and collaborative efforts to enhance the research in geothermal gradient prediction using machine learning techniques.